OpenStudy (anonymous):

given sin theta= -1/3 and sec theta > 0, find cos theta and tan theta

OpenStudy (amistre64):

sin^2 + cos^2 = 1 1/9 + cos^2 = 1 cos^2 = 1 - 1/9 cos = sqrt(8/9) = 2sqrt(2)/9

OpenStudy (amistre64):

tan = sin/cos

OpenStudy (amistre64):

since sin is - thats either in q3 or q4

OpenStudy (amistre64):

sec is cos is + in q4

OpenStudy (amistre64):

so tan is - in q4 :)

OpenStudy (anonymous):

so what does cos theta and tan theta equal?

OpenStudy (amistre64):

i gave cos; and tan, for this case, is -sin/cos

OpenStudy (amistre64):



OpenStudy (amistre64):

tan is negative so that might be a bit misleading tan = - |sin/cos| might be a better representation

OpenStudy (anonymous):

the answer include numers, though. like cos theta = - 2 the sqaure root of 2 and tan = 1/ 2 the square root of 2

OpenStudy (amistre64):

the numbers are what you get when you actually take the time to do the calculations.

OpenStudy (amistre64):

cos was calculated for you; so take the initiative to determine the tan :)